OpenStudy (anonymous):

You invest 2000 at a rate of 4% APR for a 3 months. How much interest did you earn?

OpenStudy (anonymous):

3 months = 1/4 year rate is 4 principle is 2000 interest = PRT/100

OpenStudy (anonymous):

So it's 3x4x2000?

OpenStudy (anonymous):

simple interest = principal*time*rate/100 = 2000*4*(1/4)/100 = $20

OpenStudy (anonymous):

that makes sense. how do i convert 1/4 to a decimal for a calculator though?

OpenStudy (anonymous):

type 1 divided by 4 and its a decimal

OpenStudy (anonymous):

Oh alright. Thank you sooo much.

OpenStudy (anonymous):

np ;) Just hit the blur button it helps me out alot :D